What I's the heat Capacity of 100g of copper calorimeter
1
Expert's answer
2019-08-08T10:55:21-0400

Specific heat of copper is

"S=385\mathrm{\frac{J}{kg\,C}}"

Therefore heat capacity for m=100g=0.1kg:


"C= S\cdot m=385\cdot 0.1\mathrm{\frac{J}{C}}=38.5\mathrm{\frac{J}{C}}"